A stock had returns of 11%, 1%, 9%, 15%, and -6% for the past five years. Based on these returns, what is the approximate probability that this stock will earn at least 21% in any one given year? Use population as the assumption for the data
Answer #1

Expected returns=(11%+1%+9%+15%-6%)/5=6.0000%

Standard deviation=sqrt((11%-6%)^2+(1%-6%)^2+(9%-6%)^2+(15%-6%)^2+(-6%-6%)^2)/sqrt(5)=7.5366%

Based on these returns, what is the approximate probability that this stock will earn at least 21% in any one given year?
=Pr(X>21%)
=Pr(Z>(21%-6%)/7.5366%)
=Pr(Z>1.990287)
=1-Pr(Z<1.990287)
=0.023279